OpenStudy (anonymous):

Solve the exponential equation. Exact answers only 3e^(x-1)=2+e^(x-1)

OpenStudy (anonymous):

subtracting \(e^{x-1}\) from either side gives us what?

OpenStudy (anonymous):

wait 3+2

OpenStudy (anonymous):

or 3=2

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

:/ netta pretend the \(e^{x-1}\) is another variable like \(z\):$$3z=2+z$$how would you solve for \(z\) above?

OpenStudy (anonymous):

sub both sides by z

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

and you get \(2z=2\), yes? so \(z=1\))? well here we have \(z=e^{x-1}\) hence:$$e^{x-2}=1$$can you tell me what exponent makes a power yield \(1\)?

OpenStudy (anonymous):

is it 1

OpenStudy (anonymous):

@netta nah, \(0\):$$2^0=1\\3^0=1\\\dots$$ so we know the exponent of \(e^{x-1}\) (I made a typo above) needs to be \(0\):$$x-1=0\\x=1$$